# Alleviating Distribution Shift in Synthetic Data for Machine Translation Quality Estimation

### v1

@@ Abstract
Quality Estimation (QE) models evaluate the quality of machine translations without reference translations, serving as the reward models for the translation task. Due to the data scarcity, synthetic data generation has emerged as a promising solution. However, synthetic QE data often suffers from distribution shift, which can manifest as discrepancies between pseudo and real translations, or in pseudo labels that do not align with human preferences. To tackle this issue, we introduce ADSQE, a novel framework for alleviating distribution shift in synthetic QE data. To reduce the difference between pseudo and real translations, we employ the constrained beam search algorithm and enhance translation diversity through the use of distinct generation models. ADSQE uses references—i.e., translation supervision signals—to guide both the generation and annotation processes, enhancing the quality of token-level labels. ADSE further identifies the shortest phrase covering consecutive error tokens, mimicking human annotation behavior, to assign the final phrase-level labels. Specially, we underscore that the translation model can not annotate translations of itself accurately. Extensive experiments demonstrate that ADSQE outperforms SOTA baselines like COMET in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that could benefit reward models for other tasks.
@@ 2 Background
Model architecture. Previous works, e.g. COMET Rei et al. (2023) and MQMQE Geng et al. (2023a), usually adopt multilingual pre-training language model, typically XLM-R Conneau et al. (2020), as the backbone architecture for the QE model. The outputs from the final layer of the model are used to derive representations for each token. Word-level representations are obtained by averaging the representations of all tokens within a word. Similarly, the regression score representation is computed by averaging the representations of all target tokens. These representations are subsequently fed into linear layers to predict word-level quality labels and regression scores, respectively.
@@ 3.2 Synthetic Labels
First, we parse the synthetic translation to obtain its dependency tree. Consecutive “BAD” tokens are added to the candidate set, which serves as the initial phrase. Following this, we iteratively execute the following steps:
  • Step a. We employ the Lowest Common Ancestor (LCA) algorithm Aho et al. (1973) to find the common ancestor node of the candidate tokens. This step locates the smallest sub-tree that encompasses all errors.
  • Step b. To ensure syntactic coherence, we add all tokens along the path from the LCA into the candidate set.
  • Step c. To ensure the candidate tokens are consecutive, we add all tokens, which are located between the leftmost and rightmost candidate tokens, into the candidate set.
@@ 4.1 Experiment Setup
We incorporated top-performing baselines in our experiments: COMET Rei et al. (2023) stands out as the SOTA QE model, widely adopted in translation studies Kocmi et al. (2023). COMET enhances its generalization capabilities by leveraging labeled QE datasets with various annotations [^3] https://github.com/sheffieldnlp/mlqe-pe, across multiple language pairs. Additionally, COMET is built on the XLMR-XL model Conneau et al. (2020), which has seven times more parameters than ours. GEMBA-MQM Kocmi and Federmann (2023) employs few-shot prompts to guide GPT-4 in generating MQM predictions. InstructScore Xu et al. (2023) and MQMQE Geng et al. (2023a) are two representative synthetic data approaches. InstructScore generates MQM data by prompting GPT-4, whereas MQMQE employs a translation model combined with negative sampling to produce synthetic data.
@@
Since COMET does not provide sentence-level results for the individual model, we reproduce sentence-level results using the released model [^5] https://huggingface.co/Unbabel/wmt23-cometkiwi-da-xl. For word- and span-level results, we directly utilize the results provided in the COMET report Rei et al. (2023), as the corresponding word- and span-level models have not been released. InstructScore is implemented using 10K synthetic data released by Xu et al. (2023). For ADSQE and MQMQE, we generate 500K synthetic data in each language pair. In Section 4.5, we demonstrate that ADSQE remains competitive with InstructScore when trained on 10K synthetic data while requiring significantly lower costs. Additional implementation details are provided in Appendix A.
@@ 4.2 Main Results
As demonstrated in Table 2, ADSQE substantially outperforms COMET despite utilizing fewer parameters, achieving notable improvements with average gains of 4.38 in Spearman, 3.41 in MCC, and 1.45 in F1 score. Moreover, ADSQE significantly outperforms GEMBA-MQM, which is based on the advanced LLM, GPT-4. Compared to other synthetic methods, i.e., MQMQE and InstructScore, ADSQE demonstrates consistent superiority, indicating that our synthetic data is of higher quality.
@@
As shown in Table 2, both MQMQE and InstructScore demonstrate significant performance declines compared to their supervised counterparts, with reductions of 15.74 and 7.64 on average, respectively. That implies the distribution shift problem in previous synthetic QE data. In contrast, our proposed method achieves superior robustness, incurring a smaller average reduction of 6.64 points. Moreover, ADSQE outperforms COMET, which relies on labeled datasets from other language pairs, on HE-EN. This indicates that our synthetic data provides more relevant QE knowledge for HE-EN compared to datasets from different language pairs.
@@ 5 Related Works
COMET Rei et al. (2023) aims to transfer QE knowledge by leveraging labeled datasets across diverse annotations and languages. However, this idea remains limited by the scarcity of QE data. To address this limitation, synthetic data generation methods have emerged. (1) Negative sampling-based approaches like DirectQE Cui et al. (2021) and MQMQE Geng et al. (2023a), utilize translation models to generate synthetic errors through negative sampling, then directly regard error rate as sentence-level scores; (2) BSQE Tuan et al. (2021) and CBSQE Geng et al. (2023b), which generate synthetic translations based on search algorithms and derive labels by matching synthetic translations with references. In this work, we aim to alleviate the distribution shift problem by leveraging supervision signals and increasing model diversity.
@@ 6 Conclusion
The distribution shift of synthetic data poses a persistent challenge in the QE field. To address this challenge, we propose the ADSQE framework, which mitigates distribution shift by leveraging translation references, a form of translation supervision signal, to guide the generation of both diverse pseudo-translations and their corresponding pseudo-labels. Experiments show that ADSQE achieves SOTA results in both supervised and unsupervised settings. Furthermore, our analysis underscores some insights for synthetic data generation, which could benefit synthetic data methods for general reward models.

### v2

@@ Abstract
Quality Estimation (QE) models evaluate the quality of machine translations without reference translations, serving as the reward models for the translation task. Due to the data scarcity, synthetic data generation has emerged as a promising solution. However, synthetic QE data often suffers from distribution shift, which can manifest as discrepancies between pseudo and real translations, or in pseudo labels that do not align with human preferences. To tackle this issue, we introduce ADSQE, a novel framework for alleviating distribution shift in synthetic QE data. To reduce the difference between pseudo and real translations, we employ the constrained beam search algorithm and enhance translation diversity through the use of distinct generation models. ADSQE uses references—i.e., translation supervision signals—to guide both the generation and annotation processes, enhancing the quality of token-level labels. ADSE further identifies the shortest phrase covering consecutive error tokens, mimicking human annotation behavior, to assign the final phrase-level labels. Specially, we underscore that the translation model can not annotate translations of itself accurately. Extensive experiments demonstrate that ADSQE outperforms SOTA baselines like CometKiwi in both supervised and unsupervised settings. Further analysis offers insights into synthetic data generation that could benefit reward models for other tasks.
@@ 2 Background
Model architecture. Previous works, e.g. CometKiwi Rei et al. (2023) and MQMQE Geng et al. (2023a), usually adopt multilingual pre-training language model, typically XLM-R Conneau et al. (2020), as the backbone architecture for the QE model. The outputs from the final layer of the model are used to derive representations for each token. Word-level representations are obtained by averaging the representations of all tokens within a word. Similarly, the regression score representation is computed by averaging the representations of all target tokens. These representations are subsequently fed into linear layers to predict word-level quality labels and regression scores, respectively.
@@ 3.2 Synthetic Labels
First, we parse the synthetic translation to obtain its dependency tree. Consecutive “BAD” tokens are added to the candidate set, which serves as the initial phrase (i.e. “action with his”). Following this, we iteratively execute the following steps:
  • Step a. We employ the Lowest Common Ancestor (LCA) algorithm Aho et al. (1973) to find the common ancestor node of the candidate tokens (i.e. word “take”). This step locates the smallest sub-tree that encompasses all errors.
  • Step b. To ensure syntactic coherence, we add all tokens along the path from the LCA into the candidate set (i.e. the word “consent” is added as it lies on the path from “his” to “ take”).
  • Step c. To ensure the candidate tokens are consecutive, we add all tokens, which are located between the leftmost and rightmost candidate tokens, into the candidate set (i.e. the word “some” is added as it appears within the phrase “take some action with his consent”).
@@ 4.1 Experiment Setup
We incorporated top-performing baselines in our experiments: CometKiwi Rei et al. (2023) stands out as the SOTA QE model, widely adopted in translation studies Kocmi et al. (2023). CometKiwi enhances its generalization capabilities by leveraging labeled QE datasets with various annotations [^3] https://github.com/sheffieldnlp/mlqe-pe, across multiple language pairs. Additionally, CometKiwi is built on the XLMR-XL model Conneau et al. (2020), which has seven times more parameters than ours. GEMBA-MQM Kocmi and Federmann (2023) employs few-shot prompts to guide GPT-4 in generating MQM predictions. InstructScore Xu et al. (2023) and MQMQE Geng et al. (2023a) are two representative synthetic data approaches. InstructScore generates MQM data by prompting GPT-4, whereas MQMQE employs a translation model combined with negative sampling to produce synthetic data.
@@
Since CometKiwi does not provide sentence-level results for the individual model, we reproduce sentence-level results using the released model [^5] https://huggingface.co/Unbabel/wmt23-cometkiwi-da-xl. For word- and span-level results, we directly utilize the results provided in the CometKiwi report Rei et al. (2023), as the corresponding word- and span-level models have not been released. InstructScore is implemented using 10K synthetic data released by Xu et al. (2023). For ADSQE and MQMQE, we generate 500K synthetic data in each language pair. In Section 4.5, we demonstrate that ADSQE remains competitive with InstructScore when trained on 10K synthetic data while requiring significantly lower costs. Additional implementation details are provided in Appendix A.
@@ 4.2 Main Results
As demonstrated in Table 2, ADSQE substantially outperforms CometKiwi despite utilizing fewer parameters, achieving notable improvements with average gains of 4.38 in Spearman, 3.41 in MCC, and 1.45 in F1 score. Moreover, ADSQE significantly outperforms GEMBA-MQM, which is based on the advanced LLM, GPT-4. Compared to other synthetic methods, i.e., MQMQE and InstructScore, ADSQE demonstrates consistent superiority, indicating that our synthetic data is of higher quality.
@@
As shown in Table 2, both MQMQE and InstructScore demonstrate significant performance declines compared to their supervised counterparts, with reductions of 15.74 and 7.64 on average, respectively. That implies the distribution shift problem in previous synthetic QE data. In contrast, our proposed method achieves superior robustness, incurring a smaller average reduction of 6.64 points. Moreover, ADSQE outperforms CometKiwi, which relies on labeled datasets from other language pairs, on HE-EN. This indicates that our synthetic data provides more relevant QE knowledge for HE-EN compared to datasets from different language pairs.
@@ 5 Related Works
CometKiwi Rei et al. (2023) aims to transfer QE knowledge by leveraging labeled datasets across diverse annotations and languages. However, this idea still faces the pitfall of distribution shift, as highlighted in Zouhar et al. (2024), primarily due to the scarcity of QE data. To address this limitation, synthetic data generation methods have emerged. (1) Negative sampling-based approaches like DirectQE Cui et al. (2021) and MQMQE Geng et al. (2023a), utilize translation models to generate synthetic errors through negative sampling, then directly regard error rate as sentence-level scores; (2) BSQE Tuan et al. (2021) and CBSQE Geng et al. (2023b), which generate synthetic translations based on search algorithms and derive labels by matching synthetic translations with references. In this work, we aim to alleviate the distribution shift problem by leveraging supervision signals and increasing model diversity.
@@ 6 Conclusion
The distribution shift of synthetic data poses a persistent challenge in the QE field. To tackle this challenge, we introduce the ADSQE framework, which mitigates distribution shifts by utilizing translation references (a form of supervision signal) and enhancing the diversity among translation models. Experiments show that ADSQE achieves SOTA results in both supervised and unsupervised settings. Furthermore, our analysis underscores some insights for synthetic data generation, which could benefit synthetic data methods for general reward models.